In an introductory statistics course, let x equal the midterm exam score and y equal the final exam score. Both have a mean of 84 and a standard deviation of 6. The correlation between the exam scores is 0.68. a. Find the regression equation. b. Find the predicted final exam score for a student with a midterm score of 84 and another with a midterm exam score of 93.
Answer #1

a)here slope b1=r*Sy/Sx =0.68*6/6 =0.68

and intercept =84-0.68*84 =26.88

therefore  regression equation: Y^ =26.88+0.68x

b)

predicted final exam score for a student with a midterm score of 84 =26.88+0.68*84 =84

predicted final exam score for a student with a midterm score of 93 =26.88+0.68*93 =90.12